# Angular Momentum Drain: Despinning Embedded Planetesimals

### v1

@@ Abstract
Young and forming planetesimals experience impacts from particles present in a protostellar disk. Using crater scaling laws, we integrate ejecta distributions for oblique impacts. For impacts at 10 to 65 m/s, expected for impacts associated with a disk wind, we estimate the erosion rate and torque exerted on the planetesimal. We find that the mechanism for angular momentum drain proposed by Dobrovolskis and Burns (1984) for asteroids could operate in the low velocity regime of a disk wind. Though spin-down associated with impacts can facilitate planetesimal collapse, we find that the process is inefficient. We find that angular momentum drain via impacts operates in the gravitational focusing regime, though even less efficiently than for lower mass planetesimals. The angular momentum transfer is most effective when the wind speed is low, the projectile density is high compared to the bulk planetesimal density, and the planetesimal is composed of low-strength material. Due to its inefficiency, we find that angular momentum drain due to impacts within a pebble cloud does not by itself facilitate collapse of single planetesimals.
@@ 1 Introduction
The spin rates of most asteroids have been altered by impacts (Farinella et al., 1992; Mao and McKinnon, 2020). Because high velocity ejecta preferentially escapes along the direction of rotation, impacts cumulatively cause asteroids and other minor bodies to spin down (Dobrovolskis and Burns, 1984; Cellino et al., 1990; Takeda and Ohtsuki, 2009; Ševeček et al., 2019). Additional mechanisms for minor body spin evolution include sublimation and outgassing (Mueller and Ferrin, 1996; Mottola et al., 2020), the YORP effect (Bottke et al., 2006) and tidal encounters (Scheeres et al., 2004). In this study we focus on the mechanism for angular momentum drain by impacts proposed by Dobrovolskis and Burns (1984), but apply it in the context of recently formed planetesimals during the epoch when they are embedded in a circumstellar disk.
@@
Planetesimals are thought to form in particle-rich clumps that coalesce within a protoplanetary disk during an era of streaming instability (Johansen et al., 2007; Li et al., 2019; Carrera et al., 2021) or within turbulent vortices (Cuzzi et al., 2008). For variations or alternatives to these scenarios, see the review by Simon et al. (2022). Planetesimal collapse within a cloud of pebbles is expected to be inefficient, with only a fraction (between 50% to 100%) of a pebble cloud mass incorporated into bound objects (Robinson et al., 2020; Nesvorný et al., 2021; Polak and Klahr, 2023). During the final states of planetesimal collapse, the angular momentum of the pebble cloud can generate a rapidly spinning disk or a bound binary (Nesvorný et al., 2019; Robinson et al., 2020; Nesvorný et al., 2021; Polak and Klahr, 2023). A newly formed planetesimal can experience accretion, erosion, and transport of particles across its surface while embedded in the protoplanetary disk (Rozner et al., 2020; Demirci et al., 2020; Quillen et al., 2024).
For a clump of mass $M$ to be self-gravitating, its self-gravity should exceed the tidal force from the host star. A gas or particle cloud would collapse if its mean density $\rho_{c}\gtrsim\rho_{H*}$ where the modified Hill density is
$\rho_{H*}$ $=\frac{1}{\pi}\left(\frac{3}{2}\right)^{5}\frac{M_{*}}{a^{3}},$ (1)
$=1.4\times 10^{-3}{\rm\ kg~{}m}^{-3}\left(\frac{M_{*}}{M_{\odot}}\right)\left(\frac{a}{1\ {\rm AU}}\right)^{-3}$
(Paczyński 1971, but also following equation 2 by Polak and Klahr 2023). Here $M_{*}$ is the mass of the star and $a$ is the orbital semi-major axis about the star. At an orbital semi-major axis of $a=45$ AU with $M_{*}=M_{\odot}$, the modified Hill density is $\rho_{H*}=1.54\times 10^{-8}{\rm\ kg~{}m}^{-3}$. This is approximately $10^{11}$ times smaller than the density of Kuiper Belt Objects, ranging from $235\ {\rm kg~{}m^{-3}}$ to $2500\ {\rm kg~{}m^{-3}}$ (Brown, 2012; Bierson and Nimmo, 2019; Keane et al., 2022).
@@
If a self-gravitating pebble cloud has some spin angular momentum, during collapse it would spin up (e.g., Nesvorný et al. 2021). If the angular momentum of the cloud exceeds that of a critically rotating Jacobi ellipsoid of density about 1000 kg m-3, then angular momentum must be shed for the object to collapse into a single bound object.
@@
In the context of the formation of Transneptunian binaries, mechanisms for removing angular momentum include the ejection of a third body (Noll et al., 2008), gravitational interactions with a distribution of small bodies residing in the disk (Goldreich et al., 2002; Funato et al., 2004) or collisions (Weidenschilling, 2002). The simulations by Nesvorný et al. (2021) illustrate that the particles within the pebble cloud itself can serve as a reservoir for angular momentum, and that collisions between particles within the cloud can facilitate dissipation and angular momentum transfer. Binaries, contact binaries, and disk shaped objects may be a consequence of the angular momentum of a pebble cloud prior to its collapse (Nesvorný et al., 2021; Polak and Klahr, 2023).
@@
We focus on the connection between impacts from particles in a pebble cloud and the spin angular momentum of a collapsing body. As impacts can spin down an asteroid (Dobrovolskis and Burns, 1984), impacts could aid in collapse of a pebble cloud, and the host pebble cloud would serve as a reservoir or sink for angular momentum. While planetesimals are embedded in a disk, low velocity impacts could be frequent (Quillen et al., 2024). While impacts in the asteroid belt tend to be hypervelocity at a few km/s (Bottke et al., 1994) and can generate a spray of ejected particles (Lauretta et al., 2019), impact velocities on a planetesimal within a pebble cloud would be substantially lower, between a few to tens of m/s (Johansen et al., 2007; McKinnon et al., 2020; Quillen et al., 2024).
@@
With a high velocity impact on a small body, most ejecta escapes. But at lower impact velocities, the fraction of ejecta that escapes is more sensitive to body rotation, impact location, impact and ejecta angles, and the planetesimal’s escape velocity (Hirata et al., 2021; Hyodo and Genda, 2021; Quillen et al., 2024). Using an ejecta velocity distribution, we can determine the mass of ejecta that is above the escape velocity and so escapes the planetesimal. The work by Dobrovolskis and Burns (1984) dealt with high-velocity, erosional impacts that reduce the spin-rate of the asteroid. We extend this work to a low velocity regime to see how impacts affect the mass loss and angular momentum of a planetesimal that is embedded in a protostellar disk.
@@
The integrations by Dobrovolskis and Burns (1984) neglected the sensitivity of ejecta mass to azimuthal angle for oblique impacts, however recent studies (Raducan et al., 2022; Luo et al., 2022; Quillen and Doran, 2024) have begun to characterize the azimuthal dependence of oblique impact ejecta distributions. In this work, we use recent estimates for the azimuthal sensitivity of ejecta to impact angle to determine which ejecta escape. Dobrovolskis and Burns (1984) focused on the role of impacts on asteroids. In this setting the projectiles originate from different directions. The changes in angular momentum caused by impacts were modeled with a random walk (Dobrovolskis and Burns, 1984). In contrast, the spin axis of a planetesimal that is embedded within a circumstellar disk could be related to the direction of a headwind containing particle projectiles (Nesvorný et al., 2021; Quillen et al., 2024). Consequently in this study we consider impacts from a uniform direction that is perpendicular to the planetesimal spin axis.
@@ 2 Impacts on an Embedded Planetesimal
We consider a planetesimal that is bombarded by particles that reside in a circumstellar disk. We assume that, due to a difference between the velocity of the planetesimal and that of gas and small particles in the disk, impacts originate from a preferred direction, as shown in Figure 1. A planetesimal embedded in the disk feels a wind known as a ‘headwind’ that arises because the gas in the disk feels pressure, giving the gas a mean tangential circular velocity below that of a circular Keplerian orbit (Whipple, 1972). We refer to the particle flow as particles within a headwind, but impacts from particles could also occur within a vortex clump (Cuzzi et al., 2008) or a pebble cloud clump that has formed during streaming instability (Johansen et al., 2007).
@@ 2.1 Planetesimal Properties
Figure 1: An image indicating the inertial reference frame coordinate system. The direction of planetesimal rotation is shown with respect to the headwind. Dashed lines show trajectories of headwind particles that collide with the planetesimal. We assume that the planetesimal spin axis is perpendicular to the circumstellar disk plane.
@@ 2.4 Crater Scaling and Ejecta Distributions
The momentum transfer parameter and the angular momentum imparted to a planetesimal by an impact depends upon the integrated properties of the ejecta (e.g., Dobrovolskis and Burns 1984; Takeda and Ohtsuki 2009; Raducan et al. 2022). We first discuss the velocity and mass distributions of ejecta from normal impacts and then consider more general ejecta distributions from oblique impacts.
@@
where $g_{a}$ is the surface gravitational acceleration on the planetesimal surface, the projectile radius is $a_{pj}$, and the density of planetesimal and projectile are $\rho_{a}$ and $\rho_{pj}$, respectively. Impacts from headwind particles onto a planetesimal are likely to be in the strength regime (Quillen et al. 2024 following Scheeres et al. 2010) where crater and ejecta properties only depend upon $\pi_{3}$ and $\pi_{4}$.
We adopt the ejecta scaling laws for normal impacts by Housen and Holsapple (2011), which were derived assuming a point source approximation. The ejection velocity is a function of distance $x$ from the impact site where the ejecta crosses the surface level plane (as it was prior to impact), as illustrated in Figure 3. The ejecta velocity scaling law is
@@
$R_{cr,N}(\pi_{3},\pi_{4})=a_{pj}\left(\frac{4\pi}{3}\right)^{\frac{1}{3}}H_{2}\pi_{3}^{-\frac{\mu_{N}}{2}}\pi_{4}^{-\nu},$ (14)
with coefficient $H_{2}\sim 0.4$ that also depends upon substrate material properties (Housen and Holsapple, 2011).
@@
$\frac{M(<x,\pi_{4})}{m_{pj}}=\frac{3k_{N}}{4\pi}\pi_{4}\left(\left(\frac{x}{a_{pj}}\right)^{3}-n_{1}^{3}\right)$ (15)
and with coefficient $k_{N}\sim 0.3$ (Housen and Holsapple, 2011).
@@ 2.4.1 Ejecta distributions for oblique impacts
Following Raducan et al. (2022); Quillen and Doran (2024) we modify equations 13, 14, and 15 to be sensitive to both impact angle and azimuthal angle, giving functions for crater radius $R_{cr}$, exponent $\mu$, and coefficients $C_{1},k$ (dropping the $N$ subscript) that are consistent with normal impact scaling
@@
The functions given by Raducan et al. (2022) in their equations 6–9 (and here with our equations 2023) are not consistent with their Figure 12 but appear consistent with the distributions shown in their other figures. The weak dependence of the function $k(\zeta,\theta_{imp})$ on angles $\theta_{imp}$ and $\zeta$ is supported by laboratory measurements of the azimuthal ejecta mass distribution from 104 m/s oblique impacts into sand (Quillen and Doran, 2024).
@@
$\frac{M(<x,\pi_{4},\theta_{imp},\zeta)}{m_{pj}}=\frac{3k(\theta_{imp},\zeta)}{4\pi}\pi_{4}\left(\left(\frac{x}{a_{pj}}\right)^{3}-n_{1}^{3}\right)$ (25)
@@ 2.5 Ejecta Velocity Components
Ejecta distributions are calculated in the frame rotating with the surface; however, to determine which ejecta escape the planetesimal, we compare the ejecta velocities in the inertial frame to the escape velocity. Because the maximum ejecta velocity is usually less than the projectile velocity, the ratio of ejecta velocity to rotational velocity is larger than the ratio of projectile velocity to rotational velocity. This means when we transfer back into the inertial frame, planetesimal rotation causes a more significant change in the ejecta velocity, than there was when we transferred into the rotating frame. This is illustrated in Figures 2 - 4. In other words, we could have neglected the rotational velocity when estimating the ejecta distribution, but should not neglect the rotational velocity when determining which ejecta escapes.
In the surface rotational frame, and in terms of the surface normal, downrange and binormal vector directions, the ejecta velocity is
@@ 2.6 The Mass and Momentum of Escaping Ejecta
$\frac{d\left(M(<x,\pi_{4},\theta_{imp},\zeta)\right)}{m_{pj}}=\frac{9k(\theta_{imp},\zeta)}{4\pi}\pi_{4}x^{2}dxd\zeta.$ (30)

### v2

@@ Abstract
Young and forming planetesimals experience impacts from particles present in a protostellar disk. Using crater scaling laws, we integrate ejecta distributions for oblique impacts. For impacts at 10 to 65 m/s, expected for impacts associated with a disk wind, we estimate the erosion rate and torque exerted on the planetesimal. We find that the mechanism for angular momentum drain proposed by Dobrovolskis and Burns (1984) for asteroids could operate in the low velocity regime of a disk wind. Though spin-down associated with impacts can facilitate planetesimal collapse, we find that the process is inefficient. We find that angular momentum drain via impacts operates in the gravitational focusing regime. The angular momentum transfer is most effective when the wind speed is high, the projectile density is high compared to the bulk planetesimal density, and the planetesimal is sufficiently weak that cratering is in the gravity regime. We find that angular momentum drain due to impacts within a pebble cloud may, in those specific conditions, facilitate collapse of single planetesimals.
keywords:
Ejecta , Impact Phenomena , Collisional Processes , Planetesimals
@@ 1 Introduction
The spin rates of most asteroids have been altered by impacts (Farinella et al., 1992; Mao and McKinnon, 2020). Because high velocity ejecta preferentially escape along the direction of rotation, impacts cumulatively cause asteroids and other minor bodies to spin down (Dobrovolskis and Burns, 1984; Cellino et al., 1990; Takeda and Ohtsuki, 2009; Ševeček et al., 2019). Additional mechanisms for minor body spin evolution include sublimation and outgassing (Mueller and Ferrin, 1996; Mottola et al., 2020), the YORP effect (Bottke et al., 2006) and tidal encounters (Scheeres et al., 2004). In this study we focus on the mechanism for angular momentum drain by impacts proposed by Dobrovolskis and Burns (1984), but apply it in the context of recently formed planetesimals during the epoch when they are embedded in a circumstellar disk.
@@
Planetesimals are thought to form in particle-rich clumps that coalesce within a protoplanetary disk during an era of streaming instability (Youdin and Goodman, 2005; Johansen et al., 2007; Li et al., 2019; Carrera et al., 2021) or within turbulent vortices (Cuzzi et al., 2008). For variations or alternatives to these scenarios, see the review by Simon et al. (2022). Planetesimal collapse within a cloud of pebbles is expected to incorporate a fraction (between 50% to 100%) of a pebble cloud mass into bound objects (Robinson et al., 2020; Nesvorný et al., 2021; Polak and Klahr, 2023). During the final states of planetesimal collapse, the angular momentum of the pebble cloud can generate a rapidly spinning disk or a bound binary (Nesvorný et al., 2019; Robinson et al., 2020; Nesvorný et al., 2021; Polak and Klahr, 2023). A newly formed planetesimal can experience accretion, erosion, and transport of particles across its surface while embedded in the protoplanetary disk (Rozner et al., 2020; Demirci et al., 2020; Quillen et al., 2024).
@@
Nesvorný et al. (2021) find that most clumps originate with $l\sim 0.1$ at 45 AU, where $l$ is the scaled angular momentum, $l=L_{c}/L_{H}$, $L_{c}$ is the clump angular momentum, and $L_{H}$ is the Hill angular momentum. If a self-gravitating pebble cloud has some spin angular momentum, during collapse it would spin up (e.g., Nesvorný et al. 2021). As these clumps collapse, the angular momentum must be less than the angular momentum of a critically rotating Jacobi ellipsoid, $L^{*}_{J}=0.39\sqrt{GM^{3}R}$. Nesvorný et al. (2021) convert this into Hill units, $l^{*}_{J}=L^{*}_{J}/L_{H}\sim 0.01$ at 45 AU. This requires angular momentum to be shed for the object to collapse into a single bound object.
@@
In the context of the formation of Transneptunian binaries, mechanisms for removing angular momentum include the ejection of a third body (Noll et al., 2008), gravitational interactions with a distribution of small bodies residing in the disk (Goldreich et al., 2002; Funato et al., 2004), collisions (Weidenschilling, 2002), or nebular drag (McKinnon et al., 2020; Lyra et al., 2021). The simulations by Nesvorný et al. (2021) illustrate that the particles within the pebble cloud itself can serve as a reservoir for angular momentum, and that collisions between particles within the cloud can facilitate dissipation and angular momentum transfer. Binaries, contact binaries, and disk shaped objects may be a consequence of the angular momentum of a pebble cloud prior to its collapse (Nesvorný et al., 2021; Polak and Klahr, 2023).
@@
We focus on the connection between impacts from particles in a pebble cloud and the spin angular momentum of a collapsing body. As impacts can spin down an asteroid (Dobrovolskis and Burns, 1984), impacts could aid in collapse of a pebble cloud, and the host pebble cloud would serve as a reservoir or sink for angular momentum. While planetesimals are embedded in a disk, low velocity impacts could be frequent (Quillen et al., 2024). While impacts in the asteroid belt tend to be hypervelocity at a few km/s (Bottke et al., 1994) and can generate a spray of ejected particles (Jewitt et al., 2015; Lauretta et al., 2019), impact velocities on a planetesimal within a pebble cloud would be substantially lower, between a few to tens of m/s (Johansen et al., 2007; McKinnon et al., 2020; Quillen et al., 2024).
@@
With a high velocity ($\sim$ km/s) impact on a small body, most ejecta escapes. For nearly catastrophic impacts, the total escaping ejecta mass is sensitive to pre-impact body rotation (Leinhardt et al., 2000; Ballouz et al., 2015; Ševeček et al., 2019), impact angle, impact and escape velocities (Hyodo and Genda, 2021). For lower velocity impacts, due to the larger ratio between impact and escape velocities, we expect that the fraction of escaping ejecta would be more sensitive to impact angle and body rotation. At a few times escape velocity of the larger body or below, the fraction of ejecta that escapes is sensitive to impact velocity (Korycansky and Asphaug, 2009; Quillen et al., 2024), impact angle and body rotation (Ballouz et al., 2014). The works by Dobrovolskis and Burns (1984); Ševeček et al. (2019) focused on high-velocity, erosional impacts that reduce the spin-rate of an asteroid. We extend this work to a low velocity regime to see how a distribution of impacts affect the mass loss and angular momentum of a planetesimal that is embedded in a protostellar disk.
@@
The integrations by Dobrovolskis and Burns (1984) neglected the sensitivity of ejecta mass to azimuthal angle for oblique impacts, however recent studies (Raducan et al., 2022; Luo et al., 2022; Hirata et al., 2021; Quillen and Doran, 2024) have begun to characterize the azimuthal dependence of oblique impact ejecta distributions. In this work, we use recent estimates for the azimuthal sensitivity of ejecta to impact angle to determine which ejecta escape. Dobrovolskis and Burns (1984) focused on the role of impacts on asteroids. In this setting the projectiles originate from different directions. The changes in angular momentum caused by impacts were modeled with a random walk (Dobrovolskis and Burns, 1984). In contrast, the spin axis of a planetesimal that is embedded within a circumstellar disk could be related to the direction of a headwind containing particle projectiles (Nesvorný et al., 2021; Quillen et al., 2024). Consequently in this study we consider impacts from a uniform direction that is perpendicular to the planetesimal spin axis.
@@ 2 Impacts on an Embedded Planetesimal
We consider a planetesimal that is bombarded by particles that reside in a circumstellar disk. We assume that, due to a difference between the velocity of the planetesimal and that of gas and small particles in the disk, impacts originate from a preferred direction, as shown in Figure 1. A planetesimal embedded in the disk feels a wind known as a ‘headwind’ that arises because the gas in the disk is subject to a large-scale pressure gradient, giving the gas a mean tangential circular velocity below that of a circular Keplerian orbit (Whipple, 1972). We refer to the particle flow as particles within a headwind, but impacts from particles could also occur within a vortex clump (Cuzzi et al., 2008) or a pebble cloud clump that has formed during streaming instability (Johansen et al., 2007).
@@ 2.1 Planetesimal Properties
Figure 1: An image indicating the inertial reference frame coordinate system. The direction of planetesimal rotation is shown with respect to the headwind. Dashed lines show trajectories of headwind particles that collide with the planetesimal. We assume that the planetesimal spin axis is perpendicular to the circumstellar disk plane. The star is located in the positive y direction.
@@ 2.4 Crater Scaling and Ejecta Distributions
The momentum transfer parameter and the angular momentum imparted to a planetesimal by an impact depends upon the integrated properties of the ejecta (e.g., Dobrovolskis and Burns 1984; Takeda and Ohtsuki 2009; Hyodo and Genda 2021; Raducan et al. 2022). We first discuss the velocity and mass distributions of ejecta from normal impacts and then consider more general ejecta distributions from oblique impacts.
@@
where $g_{a}$ is the surface gravitational acceleration on the planetesimal surface, the projectile radius is $a_{pj}$, $Y_{a}$ is the strength of the planetesimal material, and the density of planetesimal and projectile are $\rho_{a}$ and $\rho_{pj}$, respectively.
The material properties we assume for strength regime craters is sand/fly ash (SFA), and for gravity regime craters is sand. These materials have been used as material analogues in order to study impact cratering into porous asteroids (Housen and Holsapple, 2003, 2011; Housen et al., 2018). We adopt the ejecta scaling laws for normal impacts by Housen and Holsapple (2011), which were derived assuming a point source approximation. While scaling laws are often developed in a higher velocity regime, we are applying these laws in a low-velocity regime as the experimental data compiled by Housen and Holsapple (2011) for ejecta scaling of normal impacts cover a wide range in impact speed, including impacts as low as 6 m/s.
The ejection velocity is a function of distance $x$ from the impact site where the ejecta crosses the surface level plane (as it was prior to impact), as illustrated in Figure 3. The ejecta velocity scaling law (Eq 14 from Housen and Holsapple (2011)) is
@@
$R_{cr,N}(\pi_{3},\pi_{4})=a_{pj}\left(\frac{4\pi}{3}\right)^{\frac{1}{3}}H_{2}\pi_{3}^{-\frac{\mu_{N}}{2}}\pi_{4}^{(1-3\nu)/3},$ (13)
with coefficient $H_{2}\sim 0.4$ that also depends upon substrate material properties (Housen and Holsapple, 2011). In the gravity regime, the crater radius for a normal impact is
$R_{cr,N}(\pi_{2},\pi_{4})=a_{pj}\left(\frac{4\pi}{3}\right)^{\frac{1}{3}}H_{1}\pi_{2}^{\frac{-\mu_{N}}{2+\mu_{N}}}\pi_{4}^{\frac{2+\mu_{n}-6\nu}{6+3\mu_{N}}},$ (14)
with coefficient $H_{1}\sim 0.59$. Crater scaling parameters and materials used for both strength and gravity regimes can be found in Table 2.
@@
$\frac{M(<x,\pi_{4})}{m_{pj}}=\frac{3k_{N}}{4\pi}\pi_{4}\left[\left(\frac{x}{a_{pj}}\right)^{3}-n_{1}^{3}\right]$ (15)
and with coefficient $k_{N}\sim 0.3$ (Housen and Holsapple, 2011). Section 3.1, below, contains additional discussion on the impact regime.
@@ 2.4.1 Ejecta distributions for oblique impacts
Following Raducan et al. (2022); Quillen and Doran (2024) we modify equations 12, 13 or 14, and 15 to be sensitive to both impact angle and azimuthal angle, giving functions for crater radius $R_{cr}$, exponent $\mu$, and coefficients $C_{1},k$ (dropping the $N$ subscript) that are consistent with normal impact scaling
@@
The functions given by Raducan et al. (2022) in their equations 6–9 (and here with our equations 2023) are not consistent with their Figure 12 but appear consistent with the distributions shown in their other figures. The weak dependence of the function $k(\zeta,\theta_{imp})$ on angles $\theta_{imp}$ and $\zeta$ is supported by laboratory measurements of the azimuthal ejecta mass distribution from $\sim 100$ m/s oblique impacts into sand (Quillen and Doran, 2024).
@@
$\frac{M(<x,\pi_{4},\theta_{imp},\zeta)}{m_{pj}}=\frac{3k(\theta_{imp},\zeta)}{4\pi}\pi_{4}\left[\left(\frac{x}{a_{pj}}\right)^{3}-n_{1}^{3}\right]$ (25)
@@ 2.5 Ejecta Velocity Components
Ejecta distributions are calculated in the frame rotating with the surface; however, to determine which ejecta escape the planetesimal, we compare the ejecta velocities in the inertial frame to the escape velocity. In the surface rotational frame, and in terms of the surface normal, downrange and binormal vector directions, the ejecta velocity is
@@ 2.6 The Mass and Momentum of Escaping Ejecta
$\frac{d\left(M(<x,\pi_{4},\theta_{imp},\zeta)\right)}{m_{pj}}=\frac{9k(\theta_{imp},\zeta)}{4\pi}\pi_{4}\left(\frac{x^{2}}{a_{pj}^{3}}\right)dxd\zeta.$ (30)